# Diffusion & Adversarial Schrödinger Bridges via Iterative Proportional Markovian Fitting

### v4

@@ Appendix E Experimental supplementary
Colored MNIST GitHub Link gushchin2023building MIT
CelebA Dataset Link liu2015deep Non-commercial research only
SB Benchmark GitHub Link gushchin2023building MIT
ASBM Code GitHub Link gushchin2024adversarial MIT
DSBM Code GitHub Link shi2023diffusion MIT
@@ E.2 SB benchmark $\text{B}\mathbb{W}_{2}^{2}\text{-UVP}$
We additionally study how well implementations of IPMF procedure starting from different starting processes map initial distribution $p_{0}$ into $p_{1}$ by measuring the metric $\text{B}\mathbb{W}_{2}^{2}\text{-UVP}$ also proposed by the authors of the benchmark (gushchin2023building). We present the results in Table 4. One can observe that DSBM initialized from different starting processes has quite close results and so is the case for ASBM experiments with $\epsilon\in\{1,10\}$, but with $\epsilon=0.1$ one can notice that ASBM starting from IPF and Identity experience a decline in $\text{B}\mathbb{W}_{2}^{2}\text{-UVP}$ metric.
@@
Table 4: Comparisons of $\text{B}\mathbb{W}_{2}^{2}\text{-UVP}\downarrow$ (%) between the ground truth static SB solution $p^{T}(x_{0},x_{1})$ and the learned solution on the SB benchmark. The best metric over is bolded. Results marked with $\dagger$ are taken from (gushchin2024adversarial) or (gushchin2023building).
@@ E.3 CelebA SDEdit starting processes description
We propose to take a basic image-to-image translation method and use it as a coupling to induce a starting process for the IPMF procedure. Such a coupling could provide the two properties mentioned above. We use SDEdit (meng2021sdedit) which requires an already trained diffusion model (SDE prior). Given an input image $x$, SDEdit first adds noise to the input and then denoises the resulting image by the SDE prior to make it closer to the target distribution of the SDE prior. Various models can be used as an SDE prior. We explore two options: trainable and train-free. As the first option, we train the DDPM (ho2020denoising) model on the CelebA $\times64$ size female only part. As the second option we take an already trained Stable Diffusion (SD) V1.5 model (rombach2022high) with text prompts conditioned on which model generates $\times512$ images similar to the CelebA female part. We then apply SDEdit with the CelebA male images as input to produce similar female images using trainable DDPM and train-free SDv1.5 approaches, we call the starting processes generated by these SDEdit induced couplings DDPM-SDEdit and SD-SDEdit. Hyperparameters of SDEdit, DDPM and SDv1.5 are provided in Appendix E.9.
@@ E.4 CelebA experiment additional quantitative study
(a) FID for DSBM with various couplings.
(b) FID for ASBM with various couplings.
(c) CMMD for DSBM with various couplings.
(d) CMMD for ASBM with various couplings.
(e) MSE($x_{0}$, $\hat{x}_{1}$)
[…]
DSBM ASBM
IMF SD SDEdit DDPM SDEdit Identity IMF DDPM SDEdit SD SDEdit Identity
@@ E.5 General experimental details
Authors of ASBM (gushchin2024adversarial) kindly provided us the code for all the experiments. All the hyperparameters including neural networks architectures were chosen as close as possible to the ones used by the authors of ASBM in their experimental section. Particularly, as it is described in (gushchin2024adversarial, Appendix D), authors used DD-GAN (xiaotackling) with Brownian Bridge posterior sampling instead of DDPM’s one and implementation from:
DSBM (shi2023diffusion) implementation is taken from the official code repository:
Sampling on the inference stage is done by Euler Maryama SDE numerical solver (kloeden1992) with indicated in Table 6 NFE.
@@ E.6 Illustrative 2D examples details
ASBM. For toy experiments the MLP with hidden layers $[256,256,256]$ has been chosen for both discriminator and generator. The generator takes vector of $(dim+1+2)$ length with data, latent variable and embedding (a simple lookup table torch.nn.Embedding) dimensions, respectively. The networks have torch.nn.LeakyReLU as activation layer with $0.2$ angle of negative slope. The optimization has been conducted using torch.optim.Adam with running averages coefficients $0.5$ and $0.9$. Additionally, the CosineAnnealingLR scheduler has been used only at pretraining iteration with minimal learning rate set to 1e-5 and no restarting. To stabilize GAN training R1 regularizer with coefficient $0.01$ (mescheder2018training) has been used.
@@ E.7 SB benchmark details
Scrödinger Bridges/Entropic Optimal Transport Benchmark (gushchin2023building) and $\text{cB}\mathbb{W}_{2}^{2}\text{-UVP}$, $\text{B}\mathbb{W}_{2}^{2}\text{-UVP}$ metric implementation was taken from the official code repository:
@@ E.8 CMNIST details
Working with the MNIST dataset, we use a regular train/test split with 60000 images and 10000 images respectively. We RGB color train and test digits of classes “2” and “3”. Each sample is resized to $32\times 32$ and normalized by $0.5$ mean and $0.5$ std.
@@
ASBM. The cMNIST setup mainly differs by the architecture used. The generator model is built upon the NCSN++ architecture (song2021sde), following the approach in (xiaotackling) and (gushchin2024adversarial). We use 2 residual and attention blocks, 128 base channels, and $(1,2,2,2)$ feature multiplications per corresponding resolution level. The dimension of the latent vector has been set to 100. Following the best practices of time-dependent neural networks sinusoidal embeddings are employed to condition on the integer time steps, with a dimensionality equal to $2\times$ the number of initial channel, resulting in a 256-dimensional embedding. The discriminator adopts ResNet-like architecture with 4 resolution levels. The same optimizer with the same parameters as in toy E.6 and SB benchmark E.7 experiments have been used except ones that are presented in Table 6. No scheduler has been applied. Additionally, R1 regularization is applied to the discriminator with a coefficient of 0.02, in line with (xiaotackling) and (gushchin2024adversarial).
DSBM. The model is based on the U-Net architecture (ronneberger2015u) with attention blocks, 2 residual blocks per level, 4 attention heads, 128 base channels, $(1,2,2,2)$ feature multiplications per resolution level. Training was held by Adam (kingma2014adam) optimizer.
@@ E.9 CelebA details
Test FID, see Figure 6 is calculated using pytorch-fid package, test CMMD is calculated using unofficial implementation in PyTorch. Working with CelabA dataset (liu2015faceattributes), we use all 84434 male and 118165 female samples ($90\%$ train, $10\%$ test of each class). Each sample is resized to $64\times 64$ and normalized by $0.5$ mean and $0.5$ std.
ASBM. As in cMNIST experiments 7 the generator model is built upon the NCSN++ architecture (song2021sde) but with small parameter changes. The number of initial channels has been lowered to 64, but the number of resolution levels has been increased with the following changes in feature multiplication, which were set to $(1,1,2,2,4)$. The discriminator also has been upgraded by growing the number of resolution levels up to 6. No other changes were proposed.
@@
SDEdit coupling. DDPM (ho2020denoising) was trained on CelebA female train part processed in the same way as for other CelebA experiments. Number of diffusion steps is equal to 1000 with linear $\beta_{t}$ noise schedule, number of training steps is equal to 1M, UNet (ronneberger2015u) was used as neural network with 78M parameters, EMA was used during training with rate 0.9999. The DDPM code was taken from the official DDIM (song2020denoising) github repository:
The SDEdit method (meng2021sdedit) for DDPM model was used with 400 steps of noising and 400 steps of denoising. The code for SDEdit method was taken from the official github repository:
The Stable Diffusion V1.5 (rombach2022high) model was taken from the Huggingface (wolf2020transformers) model hub with the tag “runwayml/stable-diffusion-v1-5”. The text prompt used is “A female celebrity from CelebA”. The SDEdit method implementation for the SDv1.5 model was taken from the Huggingface library (wolf2020transformers), i.e. “StableDiffusionImg2ImgPipeline”, with hyperparameters: strength 0.75, guidance scale 7.5, number of inference steps 50. The output of SDEdit pipeline has been downscaled from $\times512$ size to $\times64$ size using bicubic interpolation.
E.10 Computational resources

### v5

@@ Appendix E Experimental supplementary
Colored MNIST GitHub Link Gushchin et al. (2023b) MIT
CelebA Dataset Link Liu et al. (2015a) Non-commercial research only
SB Benchmark GitHub Link Gushchin et al. (2023b) MIT
ASBM Code GitHub Link Gushchin et al. (2024b) MIT
DSBM Code GitHub Link Shi et al. (2023) MIT
@@ E.2 SB benchmark $\text{B}\mathbb{W}_{2}^{2}\text{-UVP}$
We additionally study how well implementations of IPMF procedure starting from different starting processes map initial distribution $p_{0}$ into $p_{1}$ by measuring the metric $\text{B}\mathbb{W}_{2}^{2}\text{-UVP}$ also proposed by the authors of the benchmark (Gushchin et al., 2023b). We present the results in Table 5. One can observe that DSBM initialized from different starting processes has quite close results and so is the case for ASBM experiments with $\epsilon\in\{1,10\}$, but with $\epsilon=0.1$ one can notice that ASBM starting from IPF and Identity experience a decline in $\text{B}\mathbb{W}_{2}^{2}\text{-UVP}$ metric.
@@
Table 5: Comparisons of $\text{B}\mathbb{W}_{2}^{2}\text{-UVP}\downarrow$ (%) between the ground truth static SB solution $p^{T}(x_{0},x_{1})$ and the learned solution on the SB benchmark. The best metric over is bolded. Results marked with $\dagger$ are taken from (Gushchin et al., 2024b) or (Gushchin et al., 2023b).
@@ E.3 CelebA SDEdit starting processes description
We propose to take a basic image-to-image translation method and use it as a coupling to induce a starting process for the IPMF procedure. Such a coupling could provide the two properties mentioned above. We use SDEdit (Meng et al., 2022) which requires an already trained diffusion model (SDE prior). Given an input image $x$, SDEdit first adds noise to the input and then denoises the resulting image by the SDE prior to make it closer to the target distribution of the SDE prior. Various models can be used as an SDE prior. We explore two options: trainable and train-free. As the first option, we train the DDPM (Ho et al., 2020) model on the CelebA $\times64$ size female only part. As the second option we take an already trained Stable Diffusion (SD) V1.5 model (Rombach et al., 2022) with text prompts conditioned on which model generates $\times512$ images similar to the CelebA female part. We then apply SDEdit with the CelebA male images as input to produce similar female images using trainable DDPM and train-free SDv1.5 approaches, we call the starting processes generated by these SDEdit induced couplings DDPM-SDEdit and SD-SDEdit. Hyperparameters of SDEdit, DDPM and SDv1.5 are provided in Appendix E.9.
@@ E.4 CelebA experiment additional quantitative study
(a) $x\sim p_{0}$}
(b) DSBM-Identity
(c) ASBM-Identity
Figure 9: Results on the CelebA dataset for the male $\rightarrow$ female translation task, where $x_{0}\sim p_{0}$} represents samples from the source distribution. DSBM-Identity and ASBM-Identity refers to the outputs generated using trained DSBM/ASBM with the Identity initialization. The model was trained with $\epsilon=10$.
Initialisation (coupling) DSBM ASBM
IMF DDPM SDEdit SD SDEdit Identity IMF DDPM SDEdit SD SDEdit Identity Identity $\epsilon{=}10$ IMF DDPM SDEdit SD SDEdit Identity Identity $\epsilon{=}10$
@@
Initialisation (coupling) DSBM ASBM
IMF DDPM SDEdit SD SDEdit Identity IMF DDPM SDEdit SD SDEdit Identity IMF DDPM SDEdit SD SDEdit Identity
[…]
(a) CMMD for DSBM with various couplings.
(b) CMMD for ASBM with various couplings.
@@ E.5 General experimental details
Authors of ASBM (Gushchin et al., 2024b) kindly provided us the code for all the experiments. All the hyperparameters including neural networks architectures were chosen as close as possible to the ones used by the authors of ASBM in their experimental section. Particularly, as it is described in (Gushchin et al., 2024b, Appendix D), authors used DD-GAN (Xiao et al., ) with Brownian Bridge posterior sampling instead of DDPM’s one and implementation from:
DSBM (Shi et al., 2023) implementation is taken from the official code repository:
Sampling on the inference stage is done by Euler Maryama SDE numerical solver (Kloeden, 1992) with indicated in Table 8 NFE.
@@ E.6 Illustrative 2D examples details
ASBM. For toy experiments the MLP with hidden layers $[256,256,256]$ has been chosen for both discriminator and generator. The generator takes vector of $(dim+1+2)$ length with data, latent variable and embedding (a simple lookup table torch.nn.Embedding) dimensions, respectively. The networks have torch.nn.LeakyReLU as activation layer with $0.2$ angle of negative slope. The optimization has been conducted using torch.optim.Adam with running averages coefficients $0.5$ and $0.9$. Additionally, the CosineAnnealingLR scheduler has been used only at pretraining iteration with minimal learning rate set to 1e-5 and no restarting. To stabilize GAN training R1 regularizer with coefficient $0.01$ (Mescheder et al., 2018) has been used.
@@ E.7 SB benchmark details
Scrödinger Bridges/Entropic Optimal Transport Benchmark (Gushchin et al., 2023b) and $\text{cB}\mathbb{W}_{2}^{2}\text{-UVP}$, $\text{B}\mathbb{W}_{2}^{2}\text{-UVP}$ metric implementation was taken from the official code repository:
@@ E.8 CMNIST details
Working with the MNIST dataset, we use a regular train/test split with 60000 images and 10000 images respectively. We RGB color train and test digits of classes “2” and “3”. Each sample is resized to $32\times 32$ and normalized by $0.5$ mean and $0.5$ std. ASBM. The cMNIST setup mainly differs by the architecture used. The generator model is built upon the NCSN++ architecture (Song et al., ), following the approach in (Xiao et al., ) and (Gushchin et al., 2024b). We use 2 residual and attention blocks, 128 base channels, and $(1,2,2,2)$ feature multiplications per corresponding resolution level. The dimension of the latent vector has been set to 100. Following the best practices of time-dependent neural networks sinusoidal embeddings are employed to condition on the integer time steps, with a dimensionality equal to $2\times$ the number of initial channel, resulting in a 256-dimensional embedding. The discriminator adopts ResNet-like architecture with 4 resolution levels. The same optimizer with the same parameters as in toy E.6 and SB benchmark E.7 experiments have been used except ones that are presented in Table 8. No scheduler has been applied. Additionally, R1 regularization is applied to the discriminator with a coefficient of 0.02, in line with (Xiao et al., ) and (Gushchin et al., 2024b).
@@
DSBM. The model is based on the U-Net architecture (Ronneberger et al., 2015) with attention blocks, 2 residual blocks per level, 4 attention heads, 128 base channels, $(1,2,2,2)$ feature multiplications per resolution level. Training was held by Adam (Kingma and Ba, 2014) optimizer.
@@ E.9 CelebA details
Test FID, see Figure 5 is calculated using pytorch-fid package, test CMMD is calculated using unofficial implementation in PyTorch. Working with CelabA dataset (Liu et al., 2015b), we use all 84434 male and 118165 female samples ($90\%$ train, $10\%$ test of each class). Each sample is resized to $64\times 64$ and normalized by $0.5$ mean and $0.5$ std.
ASBM. As in cMNIST experiments E.8 the generator model is built upon the NCSN++ architecture (Song et al., ) but with small parameter changes. The number of initial channels has been lowered to 64, but the number of resolution levels has been increased with the following changes in feature multiplication, which were set to $(1,1,2,2,4)$. The discriminator also has been upgraded by growing the number of resolution levels up to 6. No other changes were proposed.
@@
SDEdit coupling. DDPM (Ho et al., 2020) was trained on CelebA female train part processed in the same way as for other CelebA experiments. Number of diffusion steps is equal to 1000 with linear $\beta_{t}$ noise schedule, number of training steps is equal to 1M, UNet (Ronneberger et al., 2015) was used as neural network with 78M parameters, EMA was used during training with rate 0.9999. The DDPM code was taken from the official DDIM (Song et al., 2021) github repository:
The SDEdit method (Meng et al., 2022) for DDPM model was used with 400 steps of noising and 400 steps of denoising. The code for SDEdit method was taken from the official github repository:
The Stable Diffusion V1.5 (Rombach et al., 2022) model was taken from the Huggingface (Wolf et al., 2020) model hub with the tag “runwayml/stable-diffusion-v1-5”. The text prompt used is “A female celebrity from CelebA”. The SDEdit method implementation for the SDv1.5 model was taken from the Huggingface library (Wolf et al., 2020), i.e. “StableDiffusionImg2ImgPipeline”, with hyperparameters: strength 0.75, guidance scale 7.5, number of inference steps 50. The output of SDEdit pipeline has been downscaled from $\times512$ size to $\times64$ size using bicubic interpolation.
E.10 AFHQ details
We first pretrain the networks using Bridge Matching for 100000 steps, then run DSBM for 20 iterations with 25000 steps per outer iteration. We follow (Shi et al., 2023) and use the same U-Net architecture. The batch size is 4, and the EMA rate is 0.999. We choose $\sigma^{2}$ = 5, and again we use 100 sampling steps with constant stepsizes.
E.11 Computational resources